Find the missing dimension in each rectangle.
12cm. width. area=276cm 2 squared, find missing dimension.
25ft. width. area=1125ft. 2 squared, find missing dimension

Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

Find the missing dimension in each rectangle.
12cm. width. area=276cm 2 squared, find missing dimension.
Divide 276 by 12 to get the length
----------------------------------------------
25ft. width. area=1125ft. 2 squared, find missing dimension
Divide 1125 by 25 to get the length
================================================
Cheers,
Stan H.

Answer by rfer(16322)

A=Lw
L=A/w
L=276/12
L=23 cm
------------
L=1125/25
L=45 ft